OpenStudy (anonymous):

y=-x^3 +5 . I need to find the x intercept(s) and the y intercept(s). Is the graph symmetric with respect to the x-axis? Is the graph symmetric with respect to the y-axis? Is the graph symmetric with respect to the origin?

OpenStudy (amoodarya):

none of them is correct look this file

OpenStudy (anonymous):

-x^3 +5=0,the x intercept:x=5^(1/3),y=0+5,the y inteyrcept:y=5

OpenStudy (amoodarya):

the graph is symmetric with respect to the point (0,5)

OpenStudy (anonymous):

So i had the x intercept as 0 and the y as 5, but your saying the x intercept is 5^(1/3) ?

OpenStudy (anonymous):

how got the x intercept : 0 ??? @confusedstudent2012

OpenStudy (anonymous):

i mean I got 5=x^3 and i didnt know how to move on from that. I got that by subtracting the 5 out on both sides

OpenStudy (anonymous):

and then i thought, if i let x = o then y would be 5, i wasnt sure how to handle the problem

OpenStudy (amoodarya):